# Finding Kissing Numbers with Game-theoretic Reinforcement Learning

### v2

@@ Acknowledgments
We are grateful to Professor Henry Cohn for his influential work on the Kissing Number Problem, which has been a source of inspiration for this research. We are also deeply grateful for his guidance and valuable insights, particularly for helping us identify the new structures we discovered and providing constructive suggestions that greatly facilitated our progress in generalized kissing configurations. We appreciate Ao Li’s assistance in program optimization and large-scale computing cluster deployment. We thank Mingzhi Wang and Juntao Dai for their assistance in coordinating GPU resources. We thank Chiyuan Wang and Xinmian Sun for the early discussions on this topic.
@@
C.M. conceived and led the project. T.T.Z. established geometric metrics and analyzed the properties of the structures. C.M. established the matrix completion game and developed the game-theoretic reinforcement learning method. M.L., T.T.Z., Z.M. and C.M. explored geometric relationships between structures of different dimensions. H.C. explored the training of geometric generative models for high-dimensional structures. C.M., H.C., and M.L. modeled the process of cosine feature simulation and designed the simulation algorithm. M.L., C.M. and T.T.Z. discovered the cosine feature of constructions via numerical simulation. C.M., H.C., T.T.Z., M.L., P.L. and Z.M. completed the formulation of matrix completion and designed the algorithm. Z.M. provided practical insights and metrics for exploring lattices and error-correcting codes, particularly for the 31D construction. C.M., P.L., M.L., T.T.Z. and H.C. developed the project code. C.M., T.T.Z., P.L. and H.C. conducted the experiments. C.M. and Y.Y. wrote the manuscript. Y.Q., Y.Y. and Y.C. provided computational resources and improved the efficiency of AI systems in parallel computation.

### v3

@@ Methods
Matrix Decompositionn and Initialization for Subsequent Games.
Once a final completed matrix $\mathbf{G}^{(M)}$ is obtained, it is iteratively decomposed into a collection of submatrices
$\mathbf{G}^{(M)}_{J,J},\quad J\subset\{1,2,\dots,M\},$ (16)
where each index set $J$ selects a subset of rows and columns from $\mathbf{G}^{(M)}$. These submatrices capture representative geometric patterns and serve as distilled structural bases. For subsequent filling games, each submatrix $\mathbf{G}^{(M)}_{J,J}$ is used as an initialization matrix, providing structured guidance and constraints for Player 1 while exploring new entries. Multiple such submatrices can be used to initialize parallel filling games, allowing the system to explore diverse regions of the high-dimensional search space simultaneously. Formally, for each submatrix $\mathbf{G}^{*}=\mathbf{G}^{(M)}_{J,J}$, Player 1 resumes filling according to policy $\pi_{1}$, and Player 2 applies corrections via $\pi_{2}$, continuing the alternating process.
@@ Acknowledgments
We are grateful to Professor Henry Cohn for his influential work on the Kissing Number Problem, which has been a source of inspiration for this research. We are also deeply grateful for his guidance and valuable insights, particularly for helping us identify the new structures we discovered and providing constructive suggestions that greatly facilitated our progress in generalized kissing configurations. We appreciate the assistance of Ao Li and Xingmeng Zhang in program optimization and computing cluster deployment. We thank Mingzhi Wang and Juntao Dai in coordinating GPU resources. We thank Chiyuan Wang and Xinmian Sun for the early discussions on this topic.
@@
C.M. conceived and led the project. T.T.Z. established geometric metrics and interpreted the properties of the structures. C.M. established the matrix completion game and developed the game-theoretic reinforcement learning method. M.L., T.T.Z., Z.M. and C.M. explored geometric relationships between structures of different dimensions. H.C. explored the training of geometric generative models for high-dimensional structures. C.M., H.C. and M.L. modeled the process of cosine feature simulation and designed the simulation algorithm. M.L., C.M. and T.T.Z. discovered the cosine feature of constructions via numerical simulation. C.M., H.C., T.T.Z., M.L., P.L., Z.M. completed the formulation of matrix completion. Z.M. provided practical insights and metrics for exploring lattices and error-correcting codes, particularly for the 31-dimensional construction. C.M., P.L., M.L., T.T.Z. and H.C. developed the project code. C.M. and P.L. optimized the efficiency of overall system. C.M., T.T.Z., P.L. and H.C. conducted the experiments. C.M. and Y.Y. wrote the manuscript. Y.Q., Y.Y., and Y.C. provided computational resources and planned directions for improving the parallel efficiency. Y.Q. and Y.C. analyzed the computational time cost of the program, optimized the operator, and ensured the stable training.
@@
Supplementary Information
Appendix A Key difference to AlphaEvolve
The key difference between PackingStar and AlphaEvolve lies in the search space. AlphaEvolve optimizes explicit coordinates in $\mathbb{R}^{n}$, which is numerically unstable and scales exponentially with dimension; increasing computational resources cannot overcome this fundamental limitation. In contrast, PackingStar operates in the Gram-matrix domain with strict geometric constraints, combined with iterative decomposition into representative substructures and cooperative game dynamics that compress and distill high-dimensional search spaces. These techniques fundamentally render exploration of extremely large and complex spaces tractable, enabling efficient large-scale sampling. This fundamental distinction explains why PackingStar scales to higher dimensions while AlphaEvolve cannot [?].
Although AlphaEvolve improved the kissing number in 11 dimensions from 592 to 593, likely by adjusting the coordinates of the known 592-sphere configuration, the result lacks clear geometric structure. Furthermore, AlphaEvolve relies on a specific 11-dimensional structure with 592 spheres, which is not jammed, making it difficult to transfer similar techniques to structures in other dimensions. In contrast, PackingStar not only sets new bounds across higher dimensions but also discovers geometrically meaningful, analyzable configurations, providing concrete objects for further mathematical study.
Appendix B Diverse constructions in different dimensions